OpenStudy (jenniferjuice):

What is the simplified form of the seventh root of x to the fifth power times the seventh root of x to the fifth power? the square root of x x to the four ninth power x x times the seventh root of x cubed

OpenStudy (anonymous):

\[ \sqrt[7]{x^5} \times \sqrt[7]{x^5} = \sqrt[7]{x^{10}} = \sqrt[7]{x^3\times x^{7}} = x\sqrt[7]{x^3} \]